# Distortions in Charged-Particle Images of Laser Direct-Drive Inertial Confinement Fusion Implosions

### v1

@@ Abstract
Energetic charged particles generated by inertial confinement fusion (ICF) implosions encode information about the spatial morphology of the hot-spot and dense fuel during the time of peak fusion reactions. The knock-on deuteron imager (KoDI) was developed at the Omega Laser Facility to image these particles in order to diagnose low-mode asymmetries in the hot-spot and dense fuel layer of cryogenic deuterium–tritium ICF implosions. However, the images collected are distorted in several ways that prevent reconstruction of the deuteron source. In this paper we describe these distortions and a series of attempts to mitigate or compensate for them. We present several potential mechanisms for the distortions, including a new model for scattering of charged particles in filamentary electric or magnetic fields surrounding the implosion. A novel particle-tracing methodology is developed and utilized to create synthetic KoDI data based on the filamentary field model that reproduces the main experimentally observed image distortions. We conclude with a discussion of the outlook for KoDI, and potential considerations for other charged-particle diagnostics.
@@ I Introduction
Fusion reactions in inertial confinement fusion (ICF) implosions generate charged particles that encode information about the implosion. Some charged particles are direct products of primary or secondary fusion reactions, while other “knock-on” particles are produced by collisions between fusion neutrons and ions in the surrounding plasma. The spectra of these particles have previously been used to diagnose the yield, areal density, ion temperature, and other implosion variables [1]. The spatial distribution of the source of both fusion products and knock-on particles encodes information about the hot-spot and the dense fuel surrounding it. Imaging these particles, especially from multiple lines of sight, would allow 3D reconstruction of the burn volume and the surrounding dense fuel, allowing the identification of low-mode asymmetries that reduce implosion performance [2, 3, 4, 5].
@@
Imaging of charged-particle fusion products ($14.7$-MeV protons from D–3He reactions) was initially performed at the Omega Laser Facility with the proton core imaging system (PCIS), which was used to diagnose the radius of the burn volume [6, 7]. This technique, however, could not be simply extended to the deuterium–tritium (DT) fueled cryogenic implosions of greatest interest for ICF, which have sufficiently high areal densities to stop the lower-energy charged particles produced by the primary DT reactions.
The knock-on deuteron imager (KoDI) [8, 9] was developed to image knock-on deuterons (KoDs) and knock-on tritons (KoTs) produced by DT cryogenic implosions on the OMEGA laser. These KoDs are divided into distinct high (9 to 12 MeV) and low (3 to 6 MeV) energy populations (Fig. 1) resulting from small (${\sim}10^{\circ}$) and large (${\sim}70^{\circ}$) angle scattering, respectively [8]. High-energy deuterons detected along a particular line-of-sight (LOS) resemble a direct neutron image along that LOS. Low-energy deuterons, which must be produced by neutrons initially moving quasi-perpendicular to the detector LOS, appear to come from the dense fuel where they scatter and therefore encode information about the fuel density profile and symmetry at the neutron bang time.
@@
During the KoDI commissioning experiments, several distortions of the penumbral deuteron images were observed, as described in Sec. II. Several possible explanations for the distortions are explored in Sec. III. The initial hypothesis [8, 9] that electric fields produced by charging of the pinhole apertures were responsible provides a reasonable fit to the penumbral images, but is inconsistent with later experiments (Sec. III.1). Instead, we propose an alternative model based on the scattering of knock-on charged particles in strong filamentary electric or magnetic fields around the imploded target (Sec. III.3). A particle tracer to generate synthetic proton radiographs and synthetic KoDI data including these filamentary fields is described in Sec. IV. The particle trajectories through these fields and the synthetic KoD images are analyzed in Sec. V. Finally, we present conclusions based on this filamentary field model for KoDI and other charged-particle diagnostics in Sec. VI.
The KoDI data presented in this paper are exclusively from cryogenic implosions. A discussion of how our results relate to previous work with warm DT implosions and direct imaging of charged-particle fusion products is included in Sec. VI.
@@ II Image Distortions
Two primary distortions are observed in the KoD images collected by KoDI. First, the penumbral images are larger than expected, which we describe as “anomalous magnification” (Sec. II.1). Second, low-energy KoD penumbral images are not round, but are “smeared” across the detector, making reconstruction of the penumbral images impossible (Sec. II.2). Nonuniformities in the umbra of the penumbral images are also observed (Sec. II.3). Since the reconstruction of penumbral images is extremely sensitive to changes in the effective instrument function, the presence of these distortions makes an accurate reconstruction of the deuteron source impossible in most cases.
@@ II.1 Anomalous Magnification
Table 1: Apparent magnification determined from the x-ray and knock-on deuteron data shown in Fig. 4 using the radius of the penumbral images and the separation between penumbral images. The fitting error (appropriate for comparing $M_{\text{r,x-ray}}$ with $M_{\text{r,KoD}}$ or $M_{\text{sep,x-ray}}$ with $M_{\text{sep,KoD}}$) is reported, followed by the fitting error combined with the 2% error in the true aperture radius and separation in brackets (appropriate for comparing $M_{\text{r}}$ with $M_{\text{sep}}$).The separation magnifications agree within error. $M_{\text{r,x-ray}}$ disagrees slightly, which may be due to error in the aperture size. $M_{\text{r,x-ray}}$ and $M_{\text{r,KoD}}$ disagree significantly, indicating anomalous magnification of the deuteron images.
@@ II.2 Smearing
Figure 5: Low-energy KoD images from different shots exhibiting (a) minimal or (b) significant smearing. The points in (b) mark the center of one of the penumbral images in successive energy ranges. (c) The displacement of the center of each penumbral image (points) from the center of the highest energy bin image (stars) shows the drift clearly. (d) The displacement as a function of deuteron energy is well fit by models for deflection by electric and magnetic fields.
@@ III Possible Explanations
Several possible explanations for these image distortions have been previously explored. The most common, electrical charging of the aperture array, is discussed in Sec. III.1, while scattering of KoD in the aperture is modeled in Sec. III.2. In Sec. III.3 we propose a new explanation that we believe better fits the accumulated data.
@@ III.1 Aperture Charging
Figure 6: Azimuthally averaged high-energy deuteron data (OMEGA shot 109675) compared to the nominal PSF and PSFs from the aperture-charging model (Sec. III.1) and the filament-scattering model (Sec. III.3). The 50% radius of the data is significantly larger than the nominal PSF due to anomalous magnification.
After the initial observations of anomalous magnification, it was proposed that the effect was due to a net-negative electric charge accumulated on the aperture prior to the arrival of the deuterons [9, 8]. This charge would produce a radial electric field within each aperture that would deflect positively charged particles away from the axis, acting as an electrostatic lens to anomalously magnify the image. The radial dependence of the magnitude of the electric field could also explain some radially symmetric nonuniformities in the umbra.
@@
Another explanation for the aperture charging posited that an electromagnetic pulse radiated from the target stalk during the implosion induces a potential in the body of KoDI, drawing a negative net charge to the apertures [8]. This model predicts a linear dependence $Q_{\text{ap}}\propto L_{\text{ap}}$, consistent with the measurements in Fig. 7(a). To mitigate this, several hardware modifications were developed to attempt to prevent the flow of current from the diagnostic body to the nose tip (Fig. 3). First, a nonconductive spacer was placed between the nose cone and the body of the diagnostic. Next, the entire nose cone (except for the cap, which remained metal) was replaced with nonconductive components (plastic and ceramic). Finally, prototype apertures were manufactured out of semiconductive (silicon) and nonconductive (alumina, Al2O3) to prevent the movement of charge onto the aperture. These apertures had identical dimensions to the tantalum apertures but were thicker (${\sim}700$ $\mu$m for silicon and alumina compared to $200$ $\mu$m for tantalum) so that they remained sufficiently opaque to x-rays. The results of these experiments, shown in Fig. 7(b), show no reduction in the anomalous magnification with any of these hardware changes.
@@ III.2 Aperture Scattering
Early work on the charged particle imaging of ICF implosions [7] considered the possibility that particles could scatter in the edges of the apertures, altering the shape of the penumbral image. To investigate the possibility, we conducted Monte Carlo charged-particle scattering simulations using a particle tracker developed as part of the PlasmaPy project [20]. A point source of $10^{5}$ 6-MeV deuterons is positioned at TCC, and a detector at the same location as in the experiments. The aperture is represented by a $2$-$\mu$m resolution grid, from which the mass density and temperature at each particle location are interpolated at each time step. The mean scattering angle is calculated for each particle using cross sections from the ENDF [21] database. Finally, a histogram of particle locations in the detector plane is made and azimuthally averaged to obtain the radial profile of the penumbral image.
@@ III.3 Filamentary Fields
Figure 10: Experimental 15-MeV proton radiographs of a spherical target implosion driven by 40 beams with a 1 ns square pulse at (a) 0.39 ns, (b) 0.88 ns, and (c) 1.37 ns. Early in the pulse (a), the filamentary fields expand outward, filling the field of view later in time (b). About 300 ps after the end of the pulse (c), the filaments have largely dissipated. A dotted line shows the initial position of the target. The large feature at the top of the radiographs is caused by a current traveling down the target stalk.
@@ IV Generating Synthetic Data with a Custom Particle Pusher
Figure 11: The synthetic KoDI simulation primarily takes place in a small box surrounding the source where fields and particles are initialized. Particles are initialized with velocity based on their energy and are stopped once they reach the edge of the simulation box, or have velocity in the negative direction. Particles are then extrapolated to the aperture plane and then down-selected based on whether they make it through the aperture. The particles are then further extrapolated to the detector to form a histogram image.
@@ IV.2 Generating Synthetic KoD Images
To investigate the impact of filamentary fields on KoDI, the proton radiography particle-pushing algorithm was modified to produce synthetic KoDI data. An aperture is introduced between the deuteron source and the detector (Fig. 11). After particles leave the field grid they are evolved to the aperture plane, and particles that miss the aperture are removed before the remaining particles are propagated to the detector plane. The deuteron source is modeled as a spherical shell with a range of velocities centered on the detector. Only one energy of deuterons is traced in a given simulation, but multiple deuteron energies from multiple simulations are combined to generate synthetic data with a range of deuteron energies comparable to the experimental images. Based on the KoD scattering probability (Fig. 1), high-energy deuterons are primarily generated close to the center of the shell facing the detector, while low-energy deuterons are generated in a ring at large angles. The initial distribution of the deuteron source in the shell is weighted to reflect this.
@@ V Analysis of Synthetic KoD Images
A KoD source image is reconstructed from each synthetic penumbral image by first applying a median filter (to reduce noise from finite particle statistics) and then using the Wiener deconvolution algorithm to deconvolve the instrument PSF. The Wiener deconvolution requires an estimate of the signal-to-noise ratio, which was chosen to be $10^{-8}$ to $10^{-9}$ for the data presented here. In all cases, the nominal PSF calculated in the no-field case is used since this is what is known for the experimental data. For penumbral images generated with no electric or magnetic fields around the source, this PSF is exact and the deconvolution reproduces the source except for some noise added by the median filter. For cases with nonzero electric or magnetic fields, the reconstructed source will be distorted.
@@ V.1 Validation
Figure 14(b) shows the penumbral image produced by the same hemi-ellipsoidal source with the field on, in which the image is significantly distorted. Since scattering in the fields has effectively altered the PSF, deconvolution of the penumbral image using the nominal (no field) PSF [Fig.14(d)] no longer reproduces the true source profile. The deconvolved image is now both larger (magnified) and distorted, although the large $P_{2}$ asymmetry is still recognizable. The SSIM, as well as the absolute difference between the field and no field images, are plotted in Figs. 14(e) and 14(f), respectively, and show the significant distortion of the field-on image.
@@ V.2 Comparison to Experimental Data
Even in low-field cases where the image offset and magnification ratio appear minimal, the fields have a significant impact on the structure of the penumbral image that prevents an accurate reconstruction of the source. In the lowest electric-field case ($-32$ nC) with 12-MeV deuterons, the magnification ratio approaches unity and the image offset is $\leq 0.3$ mm, but the penumbral image profiles are significantly distorted [Fig. 17(b)]. Figures 17(e) and 17(f) show that even in this case, the fields have significantly altered the penumbra, preventing an accurate reconstruction. As a result, the SSIM plotted for these cases in Fig. 16(c) remains $<0.8$ for all energies, even as the image offset decreases and the magnification ratio approaches unity. As shown in Fig. 17(d), this level of SSIM still results in an inaccurate reconstructed source. This suggests that valid penumbral reconstructions may not be possible even for experimental data with minimal smearing or anomalous magnification.
@@ VI Conclusions
Energetic charged particles generated by ICF implosions encode information about the spatial morphology of the hot-spot and dense fuel during the time of peak fusion reactions. The KoDI was developed to image knock-on deuterons and tritons produced by implosions on the OMEGA Laser System, but the penumbral images collected are anomalously magnified, are smeared out across the detector, and exhibit other nonuniformities that together prevent the reconstruction of the deuteron source. A model based on charging of the penumbral aperture array predicts some of the distortions, but is inconsistent with other experimental data. Another model based on the scattering of deuterons in the aperture walls or within a plasma filling the aperture is found to not match the observed distortions.
We suggest that the image distortions may be better explained by a new model that describes scattering of the KoD by filamentary electric or magnetic field structures known to exist around the implosion at stagnation. A heuristic PSF based on this model provides a better fit to experimental measurements than the charged-aperture model. Particle-tracing simulations in filamentary fields based on experimental measurements reproduce all of the image distortions observed in KoDI data.
@@
We conclude that the presence of these filamentary fields is a fundamental obstacle to charged particle imaging of ICF implosions, unless the charged particles of interest are produced long after the end of the laser drive, after the filaments have dissipated. Further experiments are necessary to establish exactly how long the filaments persist. No method of mitigating the formation of filaments for ICF-relevant target materials and laser intensities is known, but this will be the subject of future work.

### v2

@@ Abstract
Energetic charged particles generated by inertial confinement fusion (ICF) implosions encode information about the spatial morphology of the hot-spot and dense fuel during the time of peak fusion reactions. The knock-on deuteron imager (KoDI) was developed at the Omega Laser Facility to image these particles in order to diagnose low-mode asymmetries in the hot-spot and dense fuel layer of cryogenic deuterium–tritium ICF implosions. However, the images collected are distorted in several ways that prevent reconstruction of the deuteron source. In this paper we describe these distortions and a series of attempts to mitigate or compensate for them. We present several potential mechanisms for the distortions, including a new model for scattering of charged particles in filamentary electric or magnetic fields surrounding the implosion. Particle-tracing is used to create synthetic KoDI data based on the filamentary field model that reproduces the main experimentally observed image distortions. We conclude that the filamentary scattering model best matches the observed image distortions. Finally, we discuss potential impacts of filamentary fields on other charged-particle diagnostics.
@@ I Introduction
Fusion reactions in inertial confinement fusion (ICF) implosions generate charged particles that encode information about the implosion. Some charged particles are direct products of primary or secondary fusion reactions, while other “knock-on” particles are produced by collisions between fusion neutrons and ions in the surrounding plasma. The spectra of these particles have previously been used to diagnose the yield, areal density, ion temperature, and other implosion variables [1]. The spatial distribution of the source of both fusion products and knock-on particles encodes information about the hot-spot and the dense fuel surrounding it. Imaging these particles from multiple lines of sight would allow 3D reconstruction of the burn volume and the surrounding dense fuel, allowing the identification of low-mode asymmetries that reduce implosion performance [2, 3, 4, 5].
@@
Imaging of charged-particle fusion products ($14.7$-MeV protons from D–3He reactions) was initially performed at the Omega Laser Facility with the proton core imaging system (PCIS), which was used to diagnose the radius of the burn volume [6, 7]. This technique could not be extended to the deuterium–tritium (DT) fueled cryogenic implosions of greatest interest for ICF, which have sufficiently high areal densities to stop the lower-energy charged particles produced by the primary DT reactions.
The knock-on deuteron imager (KoDI) [8, 9] was developed to image knock-on deuterons (KoDs) and knock-on tritons (KoTs) as well as x-rays produced by DT cryogenic implosions on the OMEGA laser. These KoDs are divided into distinct high (9 to 12 MeV) and low (3 to 6 MeV) energy populations (Fig. 1) resulting from small (${\sim}10^{\circ}$) and large (${\sim}70^{\circ}$) angle scattering, respectively [8]. High-energy deuterons detected along a particular line-of-sight (LOS) resemble a direct neutron image along that LOS. Low-energy deuterons, which must be produced by neutrons initially moving quasi-perpendicular to the detector LOS, appear to come from the dense fuel where they scatter and therefore encode information about the fuel density profile and symmetry. The KoDI KoD/KoT, and x-ray measurements are time-integrated, but are effectively time-gated by the sharp ${\sim}80$ ps duration peak in x-ray, neutron, and subsequently KoD/KoT production around the neutron bang time.
@@
During the KoDI commissioning experiments, several distortions of the penumbral deuteron images were observed, as described in Sec. II. Several possible explanations for the distortions are explored in Sec. III. The initial hypothesis [8, 9] that electric fields produced by charging of the pinhole apertures were responsible provides a reasonable fit to the penumbral images, but is inconsistent with later experiments (Sec. III.1). Here, we propose an alternative model based on the scattering of knock-on charged particles in strong filamentary electric or magnetic fields around the imploded target (Sec. III.3). A particle tracer to generate synthetic proton radiographs and synthetic KoDI data including these filamentary fields is described in Sec. IV. The particle trajectories through these fields and the synthetic KoD images are analyzed in Sec. V. Finally, we present conclusions based on this filamentary field model for KoDI and other charged-particle diagnostics in Sec. VI.
[…]
@@ II Image Distortions
Two primary distortions are observed in the KoD images collected by KoDI. First, the penumbral images are larger than expected, which we describe as “anomalous magnification” (Sec. II.1). Second, low-energy KoD penumbral images are not round, but are “smeared” across the detector, making reconstruction of the penumbral images impossible (Sec. II.2). Nonuniformities in the umbra of the penumbral images are also observed (Sec. II.3). Since the reconstruction of penumbral images is extremely sensitive to changes in the effective instrument response function, the presence of these distortions makes an accurate reconstruction of the deuteron source impossible in most cases.
@@ II.1 Anomalous Magnification
Table 1: Apparent magnification determined from the x-ray and knock-on deuteron data shown in Fig. 4 using the radius of the penumbral images and the separation between penumbral images. The fitting error (appropriate for comparing $M_{\text{r,x-ray}}$ with $M_{\text{r,KoD}}$ or $M_{\text{sep,x-ray}}$ with $M_{\text{sep,KoD}}$) is reported, followed by the fitting error combined with the 2% error in the true aperture radius and separation in brackets (appropriate for comparing $M_{\text{r}}$ with $M_{\text{sep}}$).
@@ II.2 Smearing
Figure 5: Low-energy KoD images from different shots exhibiting (a) minimal or (b) significant smearing. The points in (b) mark the center of one of the penumbral images in successive energy ranges. (c) The displacement of the center of each penumbral image (points) from the center of the highest energy bin image (stars). (d) The displacement as a function of deuteron energy is well fit by models for deflection by electric and magnetic fields.
@@ III Possible Explanations
Several possible explanations for these image distortions have been previously explored. The most common, electrical charging of the aperture array, is discussed in Sec. III.1, while scattering of KoD in the aperture is modeled in Sec. III.2. Ultimately, we show that none of these models are satisfactory. In Sec. III.3 we propose a new explanation that we believe better fits the accumulated data.
@@ III.1 Aperture Charging
Figure 6: Azimuthally averaged high-energy deuteron data (OMEGA shot 109675) compared to the nominal PSF and PSFs from the aperture-charging model (Sec. III.1) and the filament-scattering model (Sec. III.3).
After the initial observations of anomalous magnification, it was proposed that the effect was due to a net-negative electric charge accumulated on the aperture prior to the arrival of the deuterons [9, 8]. This charge would produce a radial electric field within each aperture that would deflect positively charged particles away from the axis, acting as an electrostatic lens to magnify the image. The radial dependence of the magnitude of the electric field could also explain some radially symmetric nonuniformities in the umbra.
@@
Another explanation for the aperture charging posited that an electromagnetic pulse radiated from the target stalk during the implosion induces a potential in the body of KoDI, drawing a negative net charge to the apertures [8]. This model predicts a linear dependence $Q_{\text{ap}}\propto L_{\text{ap}}$, consistent with the measurements in Fig. 7(a). To mitigate this, several hardware modifications were developed toprevent the flow of current from the diagnostic body to the nose tip (Fig. 3). First, a nonconductive spacer was placed between the nose cone and the body of the diagnostic. Next, the entire nose cone (except for the cap, which remained metal) was replaced with nonconductive components (plastic and ceramic). Finally, prototype apertures were manufactured out of semiconductive (silicon) and nonconductive (alumina, Al2O3) to prevent the movement of charge onto the aperture. These apertures had identical dimensions to the tantalum apertures but were thicker (${\sim}700$ $\mu$m for silicon and alumina compared to $200$ $\mu$m for tantalum) so that they remained sufficiently opaque to x-rays. The results of these experiments, shown in Fig. 7(b), show no reduction in the anomalous magnification with any of these hardware changes.
@@ III.2 Aperture Scattering
Early work on the charged particle imaging of ICF implosions [7] considered the possibility that particles could scatter in the edges of the apertures, altering the shape of the penumbral image. We conducted Monte Carlo charged-particle scattering simulations using a particle tracker developed as part of the PlasmaPy project [23]. A point source of $10^{5}$ 6-MeV deuterons is positioned at TCC, and a detector at the same location as in the experiments. The aperture is represented by a $2$-$\mu$m resolution grid, from which the mass density and temperature at each particle location are interpolated at each time step. The mean scattering angle is calculated for each particle using cross sections from the ENDF [24] database. Finally, a histogram of particle locations in the detector plane is made and azimuthally averaged to obtain the radial profile of the penumbral image.
@@ III.3 Filamentary Fields
Figure 10: Experimental 15-MeV proton radiographs of a spherical target implosion driven by 40 beams with a 1-ns square-shaped pulse at (a) 0.39 ns, (b) 0.88 ns, and (c) 1.37 ns. Early in the pulse (a), the filamentary fields expand outward, filling the field of view later in time (b). About 300 ps after the end of the pulse (c), the filaments have largely dissipated. A dotted line shows the initial position of the target. The large feature at the top of the radiographs is caused by a current traveling down the target stalk.
@@ IV.2 Generating Synthetic KoD Images
To investigate the impact of filamentary fields on KoDI, the proton radiography particle-pushing algorithm was modified to produce synthetic KoDI data. An aperture is introduced between the deuteron source and the detector. After particles leave the field grid they are evolved to the aperture plane, and particles that miss the aperture are removed before the remaining particles are propagated to the detector plane. The deuteron source is modeled as a spherical shell with a range of velocities centered on the detector. Only one energy of deuterons is traced in a given simulation, but multiple deuteron energies from multiple simulations are combined to generate synthetic data with a range of deuteron energies comparable to the experimental images. Based on the KoD scattering probability (Fig. 1), high-energy deuterons are primarily generated close to the center of the shell facing the detector, while low-energy deuterons are generated in a ring at large angles. The initial distribution of the deuteron source in the shell is weighted to reflect this.
@@ V Analysis of Synthetic KoD Images
A KoD source image is reconstructed from each synthetic penumbral image by first applying a median filter (to reduce noise from finite particle statistics) and then using the Wiener deconvolution algorithm to deconvolve the instrument PSF. The Wiener deconvolution requires an estimate of the signal-to-noise ratio, which was chosen to be $10^{-8}$ to $10^{-9}$ for the data presented here. In all cases, the nominal PSF calculated in the no-field case is used since this isknown for the experimental data. For penumbral images generated with no electric or magnetic fields around the source, this PSF is exact and the deconvolution reproduces the source except for some noise added by the median filter. For cases with nonzero electric or magnetic fields, the reconstructed source will be distorted.
@@ V.1 Validation
Figure 13(b) shows the penumbral image produced by the same hemi-ellipsoidal source with the field on, in which the image is significantly distorted. Since scattering in the fields has effectively altered the PSF, deconvolution of the penumbral image using the nominal (no field) PSF [Fig.13(d)] no longer reproduces the true source profile. The deconvolved image is now both larger (magnified) and distorted, although the large $P_{2}$ asymmetry is still recognizable. The SSIM, as well as the absolute difference between the field and no field images, are plotted in Figs. 13(e) and 13(f), respectively, and show the significant distortion of the field-on image.
@@ V.2 Comparison to Experimental Data
Even in low-field cases where the image offset and magnification ratio appear minimal, the fields have a significant impact on the structure of the penumbral image that prevents an accurate reconstruction of the source. In the lowest electric-field case ($-32$ nC) with 12-MeV deuterons, the magnification ratio approaches unity and the image offset is $\leq 0.3$ mm, but the penumbral image profiles are significantly distorted [Fig. 16(b)]. Figures 16(e) and 16(f) show that even in this case, the fields have significantly altered the penumbra, preventing an accurate reconstruction. As a result, the SSIM plotted for these cases in Fig. 15(c) remains $<0.8$ for all energies, even as the image offset decreases and the magnification ratio approaches unity. As shown in Fig. 16(d), this level of SSIM still results in an inaccurate reconstructed source. This suggests that valid penumbral reconstructions may not be possible even for experimental data with minimal smearing or anomalous magnification.
@@ VI Conclusions
Energetic charged particles generated by ICF implosions encode information about the spatial morphology of the hot-spot and dense fuel during the time of peak fusion reactions. The KoDI was developed to image knock-on deuterons and tritons produced by implosions on the OMEGA Laser System, but the penumbral images collected are anomalously magnified, are smeared out across the detector, and exhibit other nonuniformities that together prevent the reconstruction of the deuteron source. A model based on charging of the penumbral aperture array predicts some of the distortions, but is inconsistent with other experimental data. Another model based on the scattering of deuterons in the aperture walls or within a plasma filling the aperture does not match the observed distortions.
We suggest that the image distortionsare better explained by scattering of the KoD by filamentary electric or magnetic field structures known to exist around the implosion at stagnation. A heuristic PSF based on this model provides a better fit to experimental measurements than the charged-aperture model. Particle-tracing simulations in filamentary fields based on experimental measurements reproduce all of the image distortions observed in KoDI data.
@@
We conclude that the presence of these filamentary fields is a fundamental obstacle to charged particle imaging of ICF implosions, unless the charged particles of interest are produced long after the end of the laser drive, after the filaments have dissipated. No method of mitigating the formation of filaments for ICF-relevant target materials and laser intensities is known. We plan to conduct further experiments utilizing proton radiography and optical probing to investigate the mechanism that forms the filaments and to observe how the filaments evolve in time.
The presence of filamentary fields in the corona should also be considered when evaluating other high energy density experiments, including other potential laser direct drive fusion approaches. For example, if charged particles in a fast ignition or laboratory astrophysics experiment pass through a coronal plasma, they may be dispersed by filamentary fields.